OpenStudy (anonymous):

The vertices of the following graph are at (-2, 0) and (2, 0) True or False! I say True

OpenStudy (anonymous):

@Michele_Laino i say True

OpenStudy (michele_laino):

that's right! It is true, since they are the vertices of an hyperbola

OpenStudy (anonymous):

@Michele_Laino i said false for this

OpenStudy (michele_laino):

here we have: c=sqrt(13), so c^2=13 furthermore, from your data, I can write: a=2, and b=3. Now the subsequent equation holds (in general): \[\Large {c^2} - {a^2} = {b^2}\]

OpenStudy (michele_laino):

substituting our numeric values, for a, b, and cwe get: \[\Large {c^2} - {a^2} = 13 - 4 = 9 = {b^2}\]

OpenStudy (michele_laino):

so, what is the answer?

OpenStudy (anonymous):

9 correct

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

so the answer is false to my question since it equals 9

OpenStudy (michele_laino):

then the answer is true, since our identity is checked!

OpenStudy (anonymous):

@Michele_Laino thanks,

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

it would be A @Michele_Laino

OpenStudy (michele_laino):

sorry, I have made an error of sign, the right equation, is: \[\Large \frac{{{y^2}}}{{{a^2}}} - \frac{{{x^2}}}{{{b^2}}} = - 1\]

OpenStudy (michele_laino):

sorry again, the right equation is: \[\Large \frac{{{x^2}}}{{{a^2}}} - \frac{{{y^2}}}{{{b^2}}} = - 1\] with a=5, and b=3

OpenStudy (michele_laino):

so, waht is the right option?

OpenStudy (anonymous):

@Michele_Laino so the answer would be A ?

OpenStudy (michele_laino):

that's right! Since if we set x=0, we get y=+3 and y=-3 and they are the points of intersection of our hyperbola with the y-axis: (0,3) and (0,-3)

OpenStudy (anonymous):

@Michele_Laino i said false since the center is 1,-3 correct?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we have a=sqrt(2), and b=2, so we have to apply this equation: \[\Large \frac{{{x^2}}}{{{a^2}}} - \frac{{{y^2}}}{{{b^2}}} = - 1\] please what equation do you get, after the substitution of a and b?

OpenStudy (michele_laino):

@audicolen

OpenStudy (anonymous):

@Michele_Laino 2 and 4

OpenStudy (anonymous):

so the answer would be a @Michele_Laino

OpenStudy (michele_laino):

ok! so we get this equation: \[\frac{{{x^2}}}{2} - \frac{{{y^2}}}{4} = - 1\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

nevertheless it is not the requested equation, since, our hyperbola is traslated (from the position whose center is located at the origin of the coordinate system) by 1 unit towards right and 3 units down, so our equation has to contain the subsequent quantity: x-1, and y-3. So, waht is the right option?

OpenStudy (michele_laino):

oops! y+3, not y-3

OpenStudy (anonymous):

@Michele_Laino so the correct equation is C

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

@Michele_Laino i said true for this

OpenStudy (michele_laino):

that's right! Your answer is correct!

OpenStudy (anonymous):

last ones

OpenStudy (michele_laino):

question #7 here we have: a=5, and b=6 furthermore, we have to apply this formula: \[\frac{{{x^2}}}{{{a^2}}} - \frac{{{y^2}}}{{{b^2}}} = - 1\] So, what is the right option

OpenStudy (michele_laino):

oops! the right equation is: \[\Large \frac{{{x^2}}}{{{a^2}}} - \frac{{{y^2}}}{{{b^2}}} = 1\]

OpenStudy (michele_laino):

with a=5, and b=6, so: a^=25, and b^2=36

OpenStudy (anonymous):

@Michele_Laino so the anser would be D

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

question #8 here we have: a=4, and c=5 so we have to apply this formula: \[{c^2} - {a^2} = {b^2}\]

OpenStudy (michele_laino):

we have: \[{b^2} = 25 - 16 = ...?\] what is b^2 ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

@Michele_Laino so the correct answer is A

OpenStudy (michele_laino):

ok! so what is your option, keep in mind that we have to apply the same formula as before: \[\frac{{{x^2}}}{{{a^2}}} - \frac{{{y^2}}}{{{b^2}}} = 1\]

OpenStudy (michele_laino):

That's right! Your answer is option A

OpenStudy (michele_laino):

question #9

OpenStudy (anonymous):

number 9 is B i believe

OpenStudy (michele_laino):

we have a=3, and b=5, furthermore, we have to apply the same formula, as before: \[\frac{{{x^2}}}{{{a^2}}} - \frac{{{y^2}}}{{{b^2}}} = 1\] so, what is the right option?

OpenStudy (anonymous):

@Michele_Laino its B

OpenStudy (michele_laino):

that's right! The right option is option B

OpenStudy (anonymous):

@Michele_Laino The symmetry of a hyperbola with a center at (h, k) only occurs at y = k true or false. i said false

OpenStudy (michele_laino):

if the axis of our hyperbola is parallel to the x-axis, then the answer is true